OpenStudy (jessie13):

Will medal!!!!!!Answer ASAP!!!!! x - 2y = 3 5x + 3y = 2 The lines whose equations are shown intersect at which point? (1, -1) (-1, 1) (0, 3/2)

OpenStudy (harman.singh):

Hey jessie, those are two simultaneous equations and you can solve them by susbstitution or elimination

OpenStudy (jessie13):

???

OpenStudy (harman.singh):

Have you learnt about substitution yet to solve simultaneous equations?

OpenStudy (jessie13):

nope!!!

OpenStudy (harman.singh):

To solve a set of simultaneous equation, we need to solve one of the equations for one of the variables and substitute it into the other equation. That can help us find the P.o.I For example 2x – 3y = –2 4x + y = 24 To solve this by substitution, we rearrange one of the equations to isolate y. Lets do that with the second equation 4x+y=24 After rearranging it for y, it becomes y=24-4x . Now, we substitute this equation into our first one. Subbing y=24-4x into 2x-3y =-2 > 2x-3(24-4x)=-2 From here on, it is basic algebra and the answer will give you a value for the x coordinate. To find the the y coordinate, you simply substitute the value of x you just found into any of the two original equations. That will give you your P.o.I in the form of (x,y) If you follow this example, you should be able to figure out the answer to your confusion. Let me know if any of this didnt make sense

OpenStudy (blazeryder):

You could also use an augmented matrix. Have you hear of it before? @jessie13